OpenStudy (anonymous):

I need help with this logarithmic equation: 2log(t-2)=log(t+1) * log(t-3)

OpenStudy (anonymous):

nvm,It's actually log(t+1) + log(t-3),not * ,I can solve that

jhonyy9 (jhonyy9):

2log(t-2)=log(t-2) squared

jhonyy9 (jhonyy9):

i think that will be t2 -4t +4 = t2 -2t -3 7=2t t=7/2

jhonyy9 (jhonyy9):

because log(t+1)+log(t-3)=log((t+1)(t-3))

jhonyy9 (jhonyy9):

h

jhonyy9 (jhonyy9):

hope that is understandably and that is right

OpenStudy (anonymous):

yes,it's correct,t=7/2